OpenStudy (spareb665):

A person plans to use 90 feet of fencing and the side of his house to enclose a rectangular garden. What dimensions of the rectangle would give the maximum area?

OpenStudy (campbell_st):

so you let the sides of the rectangle be x and y so  so the perimeter is 90 = 2x + y the area is A = xy rewrite the perimeter with y as the subject 90 - 2x = y now substitute into the area formula A = x(90 - 2x) or A = -2x^2 + 90x the parabola is concave down so you will have a maximum find the 1st derivative dA/dx solve for x this will give the value of x that gives the maximum area. to find y substitute x into the perimeter formula . hope this helps

OpenStudy (campbell_st):

well if you haven't done calculus use the line of symmerty the equation is \[A = -2x^2 + 90\] for the line of symmetry use \[x = \frac{-b}{2a} \] b = 90 and a = -2 this the the length, you need to find the width by substituting into the perimeter equation.